# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

### v1

@@ 1 Introduction
The rest of this paper is organized as follows. In section 2, we provide an overview of the notations that appeared throughout the paper and present the preliminary concepts and properties of mixing matrix, conservative field, and differential inclusion. In section 3, we revisit our proposed framework (DSM) for decentralized stochastic subgradient methods, and establish the global convergence by linking it with the dynamics of its corresponding differential inclusion. In Section 4 we demonstrate how several decentralized stochastic subgradient methods can be unified within our framework, and develop the convergence properties for these methods. We present numerical experiments of analyzed decentralized subgradient methods in decentralized training nonsmooth neural networks in Section 5. In the last section, We draw conclusions and discuss some possible future research directions.
@@ 2.1 Notations
$f({\bm{x}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ (2.1)
and $F({\bm{X}}):=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}}_{i})$ with ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{d}]$. For any positive sequence $\left\{\eta_{k}\right\}$, let $\lambda(0):=0,\lambda(i):=\sum_{k=0}^{i-1}\eta_{k}$, and $\Lambda(t):=\sup\left\{k\in\mathbb{N}:t\geq\lambda(k)\right\}.$ More explicitly, $\Lambda(t)=p$, if $\lambda(p)\leq t<\lambda(p+1)$. Furthermore, we define the set-valued mapping sign : $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as follows:
$(\operatorname{sign}({\bm{x}}))_{i}=\begin{cases}\{-1\}&{\bm{x}}_{i}<0\\ {[-1,1]}&{\bm{x}}_{i}=0\\ \{1\}&{\bm{x}}_{i}>0\end{cases}$
Then it is easy to verify that $\operatorname{sign}({\bm{x}})$ is the Clarke subdifferential of the nonsmooth functions ${\bm{x}}\mapsto\|{\bm{x}}\|_{1}$. In addition, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{e}}:=\frac{1}{d}{\bm{1}}_{d}$, where ${\bm{1}}_{d}$ represents all-ones vector. Furthermore, we denote $(\Omega,\mathcal{F},\mathbb{P})$ as the probability space. We say $\left\{\mathcal{F}_{k}\right\}_{k\in\mathbb{N}}$ is a filtration if $\left\{\mathcal{F}_{k}\right\}$ is a collection of $\sigma$-algebras that satisfies $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{\infty}\subseteq\mathcal{F}$. A stochastic series $\{\xi_{k}\}$ is called a martingale with respect to a filtration $\{\mathcal{F}_{k}\}$, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=\xi_{k}$, $\forall k\in\mathbb{N}$. Moreover, A stochastic series $\{\xi_{k}\}$ is referred to as a martingale difference sequence, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=0$ holds for $\forall k\in\mathbb{N}$.
@@ 2.2 Mixing matrix
In this subsection, we introduce the concept of mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$, which determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. In general, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature.
@@
Given a connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, we say ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix of $\mathtt{G}$, if it satisfies the following conditions.
@@
With a given graph $\mathtt{G}$, there are various approaches to choose its corresponding mixing matrix, such as Laplacian-based constant edge weight matrix [43] and Metropolis weight matrix [44]. We refer interested readers to [29, 37] for more details about how to choose the mixing matrix. Based on the Definition 2.1, we present the following proposition to illustrate the fundamental properties of the mixing matrix. The proof of Proposition 2.2 follows from direct calculations, hence is omitted for simplicity.
Proposition 2.2.
For any mixing matrix ${\bm{W}}\in\mathbb{R}^{d\times d}$ that corresponds to a connected graph $\mathtt{G}$, all the eigenvalues of ${\bm{W}}$ lie in $(-1,1]$, and ${\bm{W}}$ has a single eigenvalue at $1$ that admits ${\bm{e}}$ as its eigenvector.
@@ 2.3 Nonsmooth analysis and conservative field
In this subsection, we present a brief introduction to some basic concepts in nonsmooth analysis and the conservative field [8]. A set-valued mapping $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is a mapping from $\mathbb{R}^{n}$ to set of subsets of $\mathbb{R}^{m}$. The graph of $\mathcal{G}$ is given by
$\operatorname{graph}\mathcal{G}=\{({\bm{x}},{\bm{z}})\mid{\bm{x}}\in\mathbb{R}^{n},{\bm{z}}\in\mathcal{G}({\bm{x}})\}.$
The set-valued mapping is said to be graph-closed (or has closed graph), if $\operatorname{graph}\mathcal{G}$ is a closed subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. $\mathcal{G}$ is said to be locally bounded, if for any ${\bm{x}}\in\mathbb{R}^{n}$, there exists a neighborhood $U_{{\bm{x}}}$ of ${\bm{x}}$ such that
$\sup\limits_{{\bm{z}}\in\mathcal{G}({\bm{y}}),{\bm{y}}\in U_{{\bm{x}}}}\|{\bm{z}}\|<+\infty.$
In addition, $\mathcal{G}$ is convex-valued, if $\mathcal{G}({\bm{x}})$ is a convex subset of $\mathbb{R}^{n}$ for any ${\bm{x}}\in\mathbb{R}^{n}$. Moreover, A curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is absolutely continuous if $\gamma$ is differentiable almost everywhere, and $\gamma(t)=\gamma(0)+\int_{0}^{t}\gamma^{\prime}(u)\mathrm{d}u$ holds for all $t\in\mathbb{R}_{+}$. For any locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, Rademacher’s theorem states that a locally Lipschitz continuous function $f$ is differentiable almost everywhere. We denote the subset in $\mathbb{R}^{n}$ on which $f$ is differentiable by $\Omega_{f}$. According to [13, Theorem 2.5.1], we have
@@
Then it is easy to verify that $\partial f$ is a convex-valued graph-closed locally bounded set-valued mapping from $\mathbb{R}^{n}$ to the subsets of $\mathbb{R}^{n}$. Now we present a brief introduction to the conservative field.
@@
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a nonempty set-valued mapping. We call $\mathcal{G}$ a conservative field if it is compact-valued and graph closed, and for any absolutely continuous loop $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$ satisfying $\gamma(0)=\gamma(1)$, it holds that
$\int_{0}^{1}\max_{{\bm{v}}\in\mathcal{G}(\gamma(t))}\langle\dot{\gamma}(t),{\bm{v}}\rangle\mathrm{d}t=0.$ (2.2)
We now introduce the definition of the potential function corresponding to the conservative field.
@@
Let $\mathcal{G}$ be a conservative field, we say a function $f$ is a potential for $\mathcal{G}$, provided that $f$ can be formulated as
$f({\bm{x}})={f({\bm{x}}_{0})}+\int_{0}^{t}\langle\dot{\gamma}(s),\mathcal{G}(\gamma(s))\rangle\mathrm{d}s,$
for any absolutely continuous curve $\gamma$ with $\gamma(0)={\bm{x}}_{0}$ and $\gamma(t)={\bm{x}}$. We can also say that $\mathcal{G}$ is a conservative field for $f$. Such functions $f$ are called path-differentiable functions.
As illustrated in [8, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. At the last of this part, we provide some explanations about the path-differentiable functions. In fact, the class of path-differentiable functions is general enough to cover the most objective functions in real-world problems. The well-known Clarke regular functions mentioned in [13] and semi-algebraic functions mentioned in [54] are all path-differentiable functions. In [16], the authors discuss an important function class of path-differentiable functions, definable functions, which are referred to as the functions whose graphs are definable in an o-minimal structure. Actually, most activation functions and loss functions in deep neural networks are definable functions, including sigmoid, softplus, ReLU, leaky ReLU, hinge loss, etc. Furthermore, it is important to highlight that definability remains preserved under finite summation and composition [15]. Consequently, for any neural network constructed by definable blocks, its objective function is definable, thus it is a path-differentiable function. Additionally, it is worth noting that the Clarke subdifferential of definable functions is itself definable [6]. Hence, in the case of a neural network constructed by definable blocks, the conservative field corresponding to the automatic differentiation algorithms is also definable. The following proposition demonstrates that the definability of both $f$ and its conservative field $\mathcal{G}$ leads to the nonsmooth Morse–Sard property [5].
Proposition 2.5 (Theorem 5 in [6]).
Let $f$ be a path-differentiable function that admits $\mathcal{D}_{f}$ as its conservative field. Suppose both $f$ and $\mathcal{D}_{f}$ are definable over $\mathbb{R}^{n}$, then the set $\left\{{f({\bm{x}})}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\right\}$ is finite.
Finally, based on the concept of the conservative field, we present the definition of the critical points for the optimization problem (DOP),
@@
Let $f$ in (2.1) be a path-differentiable function that admits $\mathcal{D}_{f}$ as its convex-valued conservative field, then we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a $\mathcal{D}_{f}$-critical point of (DOP), if ${\bm{X}}={\bm{X}}{\bm{e}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in\mathcal{D}_{f}({\bm{X}}{\bm{e}})$.
@@ 2.4 Differential inclusion and stochastic approximation
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a set-valued mapping. An absolutely continuous curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is called a solution (or trajectory) to the differential inclusion
$\frac{d{\bm{x}}}{dt}\in\mathcal{G}({\bm{x}}),$ (2.3)
(or a trajectory of $\mathcal{G}$) with initial point ${\bm{x}}$, provided that $\gamma(0)={\bm{x}}$ and $\gamma^{\prime}(t)\in\mathcal{G}(\gamma(t))$ holds for almost every $t\in\mathbb{R}_{+}$.
@@
Let $\mathcal{B}\subset\mathbb{R}^{n}$ be a closed set. A continuous function ϕ:nnormal-:italic-$\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is referred to as a Lyapunov function for the differential inclusion (2.3) with the stable set $\mathcal{B}$, if for any solution $\gamma$ to (2.3) and any $t>0$, it holds that
$\phi(\gamma(t))\leq\phi(\gamma(0)).$
@@
For any set-valued mapping $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and any $\delta\geq 0$, we denote $\mathcal{G}^{\delta}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as
$\mathcal{G}^{\delta}({\bm{x}})=\cup_{{\bm{y}}\in\mathbb{B}_{\delta}({\bm{x}})}(\mathcal{G}({\bm{y}})+\mathbb{B}(\delta,{\bm{0}})).$
@@
  • For above locally integrable $w$, there exists $\delta:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}\delta(t)=0$ and
    $\dot{\gamma}(t)-w(t)\in\mathcal{G}^{\delta(t)}(\gamma(t)).$
@@
where $\{\eta_{k}\}$ is a diminishing positive sequence of step-sizes, ${\bm{y}}_{k}$ is an approximate estimate of $\mathcal{G}$ at ${\bm{x}}_{k}$, and $\xi_{k+1}$ is a random noise when evaluating ${\bm{y}}_{k}$. A continuous time interpolated process $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ can be induced by (2.4),
$u(\lambda(k)+s)={\bm{x}}_{k}+\frac{{\bm{x}}_{k+1}-{\bm{x}}_{k}}{\eta_{k}}s,\quad s\in[0,\eta_{k}).$ (2.5)
Here, $\lambda(0):=0$ and $\lambda(n):=\sum_{i=0}^{n-1}\eta_{i},n\geq 1$. The following lemma shows a sufficient condition under which the interpolated process is a perturbed solution of (2.3).
Lemma 2.11 (Lemma 2.20 in [46]).
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded convex-valued graph-closed set-valued mapping. Suppose the following conditions hold in (2.4):
  1. 1.
    For any $T>0$, it holds that
    $\lim\limits_{m\rightarrow+\infty}\sup\limits_{m\leq s\leq\Lambda(\lambda(m)+T)}\left\|\sum_{k=m}^{s}\eta_{k}\xi_{k+1}\right\|=0.$
  2. 2.
    There exists a diminishing nonnegative sequence $\{\delta_{k}\}$ such that $d_{k}\in\mathcal{G}^{\delta_{k}}({x_{k}})$ holds for all $k\geq 0$.
  3. 3.
    It holds that $\sup_{k\geq 0}\left\|{\bm{x}}_{k}\right\|<+\infty$, and $\sup_{k\geq 0}\left\|{\bm{y}}_{k}\right\|<+\infty$.
Then the interpolated process $u(t)$ of $\{{\bm{x}}_{k}\}$ is a perturbed solution of (2.3).
@@
Let $\mathcal{G}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded set-valued mapping that is nonempty compact convex valued with closed graph. For any sequence $\{{\bm{x}}_{k}\}$, suppose there exists a continuous function Ψ:nnormal-:normal-$\Psi:\mathbb{R}^{n}\to\mathbb{R}$ and a closed subset $\mathcal{A}$ of $\mathbb{R}^{n}$ such that
@@
3 A General Framework for Nonsmooth Decentralized Subgradient Methods

### v2

@@ 1.4 Organizations
The rest of this paper is organized as follows. In section 2, we present some basic notations and definitions. In section 3, we describe our proposed framework (DSM) in detail and establish the global convergence by linking it with the dynamics of its corresponding differential inclusion. In Section 4, we demonstrate that framework (DSM) covers the major existing decentralized stochastic subgradient-based methods. Numerical experiments and results of analyzed decentralized subgradient-based methods are described in Section 5. In the last section, we draw conclusions and discuss possible future research directions.
@@ 2.1 Notations
For any positive sequence $\left\{\eta_{k}\right\}$, let $\lambda_{\eta}(0):=0,\lambda_{\eta}(i):=\sum_{k=0}^{i-1}\eta_{k}$, and $\Lambda_{\eta}(t):=\sup\left\{k\in\mathbb{N}:t\geq\lambda_{\eta}(k)\right\}.$ More explicitly, $\Lambda_{\eta}(t)=p$, if $\lambda_{\eta}(p)\leq t<\lambda_{\eta}(p+1)$. Furthermore, we define the set-valued mapping sign : $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as follows:
$(\operatorname{sign}({\bm{x}}))_{i}=\begin{cases}\{-1\},&{\bm{x}}_{i}<0;\\ {[-1,1],}&{\bm{x}}_{i}=0;\\ \{1\},&{\bm{x}}_{i}>0.\end{cases}$
Then it is easy to verify that $\operatorname{sign}({\bm{x}})$ is the Clarke subdifferential of the nonsmooth functions ${\bm{x}}\mapsto\|{\bm{x}}\|_{1}$.
In addition, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{1}}_{d}$ represent all-ones vector.
Furthermore, we denote $(\Omega,\mathcal{F},\mathbb{P})$ as the probability space. We say $\left\{\mathcal{F}_{k}\right\}_{k\in\mathbb{N}}$ is a filtration if $\left\{\mathcal{F}_{k}\right\}$ is a collection of $\sigma$-algebras that satisfies $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{\infty}\subseteq\mathcal{F}$. A stochastic series $\{\xi_{k}\}$ is called a martingale with respect to a filtration $\{\mathcal{F}_{k}\}$, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=\xi_{k}$, $\forall k\in\mathbb{N}$. Moreover, A stochastic series $\{\xi_{k}\}$ is referred to as a martingale difference sequence, if $\{\xi_{k}\}$ is adapted to the filtration $\{\mathcal{F}_{k}\}$ and $\mathbb{E}[\xi_{k+1}|\mathcal{F}_{k}]=0$ holds for all $k\in\mathbb{N}$.
@@ 2.2 Mixing matrix
In this subsection, we introduce the concept of mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$, which determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. Generally, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature.
@@
[44, Section 1] Given a connected graph $\mathtt{G}=(\mathtt{V},\mathtt{E})$, we say ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix of $\mathtt{G}$, if it satisfies the following conditions.
@@
With a given graph $\mathtt{G}$, there are various approaches to choose its corresponding mixing matrix, such as Laplacian-based constant edge weight matrix [45] and Metropolis weight matrix [46]. We refer interested readers to [47, 48] for more details about how to choose the mixing matrix.
Invoking [49, Perron-Frobenius Theorem], we derive a direct corollary, which illustrates the fundamental spectral property of mixing matrix ${\bm{W}}$.
Corollary 2.2.
For any mixing matrix ${\bm{W}}\in\mathbb{R}^{d\times d}$ that corresponds to a connected graph $\mathtt{G}$, all the eigenvalues of ${\bm{W}}$ lie in $(-1,1]$, and ${\bm{W}}$ has a single eigenvalue at $1$ that admits ${\bm{1}_{d}}$ as its eigenvector.
@@ 2.3 Nonsmooth analysis and conservative field
In this subsection, we present a brief introduction to some basic concepts in nonsmooth analysis and the conservative field [28].
A set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is a mapping from $\mathbb{R}^{n}$ to set of subsets of $\mathbb{R}^{m}$. The graph of $D$ is given by $\operatorname{graph}D=\{({\bm{x}},{\bm{z}})\mid{\bm{x}}\in\mathbb{R}^{n},{\bm{z}}\in D({\bm{x}})\}$. The set-valued mapping is said to be graph-closed (or has closed graph), if $\operatorname{graph}D$ is a closed subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. $D$ is said to be locally bounded, if for any ${\bm{x}}\in\mathbb{R}^{n}$, there exists a neighborhood $U_{{\bm{x}}}$ of ${\bm{x}}$ such that $\sup_{{\bm{z}}\in D({\bm{y}}),{\bm{y}}\in U_{{\bm{x}}}}\|{\bm{z}}\|<+\infty$. In addition, $D$ is convex-valued, if $D({\bm{x}})$ is a convex subset of $\mathbb{R}^{n}$ for any ${\bm{x}}\in\mathbb{R}^{n}$.
Moreover, A curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is absolutely continuous if $\gamma$ is differentiable almost everywhere, and $\gamma(t)=\gamma(0)+\int_{0}^{t}\gamma^{\prime}(u)\mathrm{d}u$ holds for all $t\in\mathbb{R}_{+}$.
For any locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, Rademacher’s theorem states that a locally Lipschitz continuous function $f$ is differentiable almost everywhere. We denote the subset in $\mathbb{R}^{n}$ on which $f$ is differentiable by $\Omega_{f}$. According to [50, Theorem 2.5.1], we have
@@
Then it is easy to verify that $\partial f$ is a convex-valued graph-closed locally bounded set-valued mapping from $\mathbb{R}^{n}$ to the subsets of $\mathbb{R}^{n}$.
Now we present a brief introduction to the conservative field.
@@
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a nonempty set-valued mapping. We call $D$ a conservative field if it is compact-valued and graph closed, and for any absolutely continuous loop $\gamma:[0,1]\rightarrow\mathbb{R}^{n}$ satisfying $\gamma(0)=\gamma(1)$, it holds that
$\int_{0}^{1}\max_{{\bm{v}}\in D(\gamma(t))}\langle\dot{\gamma}(t),{\bm{v}}\rangle\mathrm{d}t=0.$ (2.1)
It is remarkable to note that any conservative field is locally bounded [28, Remark 3]. We now introduce the definition of the potential function corresponding to the conservative field.
@@
Let $D$ be a conservative field, we say a function $f$ is a potential for $D$, provided that $f$ can be formulated as
$f({\bm{x}})={f({\bm{x}}_{0})}+\int_{0}^{t}\langle\dot{\gamma}(s),D(\gamma(s))\rangle\mathrm{d}s,$
for any absolutely continuous curve $\gamma$ with $\gamma(0)={\bm{x}}_{0}$ and $\gamma(t)={\bm{x}}$. We also say that $D$ is a conservative field for $f$, denoted as $D_{f}$. Such functions $f$ are called path-differentiable functions.
As illustrated in [28, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. More importantly, when the value of conservative field $D_{f}({\bm{x}})$ is a convex subset of $\mathbb{R}^{m}$, ${\bm{0}}\in\partial f({\bm{x}})$ implies that $0\in D_{f}({\bm{x}})$.
Finally, based on the concept of the conservative field, we present the definition of the critical points for the optimization problem (DOP).
@@
Let $f$ in (1.1) be a path-differentiable function that admits $\mathcal{D}_{f}$ as its convex-valued conservative field, then we say ${\bm{X}}\in\mathbb{R}^{n\times d}$ is a $\mathcal{D}_{f}$-critical point of (DOP), if ${\bm{X}}=\frac{1}{d}{\bm{X}}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{0}}\in\mathcal{D}_{f}(\frac{1}{d}{\bm{X}}{\bm{1}_{d}})$.
@@ 2.4 Differential inclusion and stochastic approximation
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a set-valued mapping. An absolutely continuous curve $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ is called a solution (or trajectory) to the differential inclusion
$\frac{d{\bm{x}}}{dt}\in D({\bm{x}}),$ (2.2)
(or a trajectory of $D$) with initial point ${\bm{x}}$, provided that $\gamma(0)={\bm{x}}$ and $\gamma^{\prime}(t)\in D(\gamma(t))$ holds for almost every $t\in\mathbb{R}_{+}$.
@@
Let $\mathcal{B}\subset\mathbb{R}^{n}$ be a closed set. A continuous function $\psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is referred to as a Lyapunov function for the differential inclusion (2.2) with the stable set $\mathcal{B}$, if for any solution $\gamma$ to (2.2) and any $t>0$, it holds that
$\psi(\gamma(t))\leq\psi(\gamma(0)).$
@@
For any set-valued mapping $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and any $\delta\geq 0$, we denote $D^{\delta}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ as $D^{\delta}({\bm{x}})=\cup_{{\bm{y}}\in\mathbb{B}_{\delta}({\bm{x}})}(D({\bm{y}})+\mathbb{B}(\delta,{\bm{0}}))$.
@@
  • For above locally integrable $w$, there exists $\delta:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}\delta(t)=0$ and $\dot{\gamma}(t)-w(t)\in D^{\delta(t)}(\gamma(t))$.
@@
where $\{\eta_{k}\}$ is a diminishing positive sequence of step-sizes, ${\bm{y}}_{k}$ is an approximate estimate of $D$ at ${\bm{x}}_{k}$, and $\xi_{k+1}$ is a random noise when evaluating ${\bm{y}}_{k}$. A continuous time interpolated process $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ can be induced by (2.3),
$u(\lambda_{\eta}(k)+s)={\bm{x}}_{k}+\frac{{\bm{x}}_{k+1}-{\bm{x}}_{k}}{\eta_{k}}s,\quad s\in[0,\eta_{k}).$ (2.4)
Here, $\lambda_{\eta}(0):=0$ and $\lambda_{\eta}(n):=\sum_{i=0}^{n-1}\eta_{i},n\geq 1$.
@@
Let $D:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ be a locally bounded set-valued mapping that is nonempty compact convex valued with closed graph. For any sequence $\{{\bm{x}}_{k}\}$, suppose there exists a continuous function $\psi:\mathbb{R}^{n}\to\mathbb{R}$ and a closed subset $\mathcal{A}$ of $\mathbb{R}^{n}$ such that
@@
3 A General Framework for Decentralized Stochastic Subgradient-based Methods